Please help fan+medal!! Find the probability of choosing the same activity when two friends each randomly choose 1 of 4 extracurricular activities to participate in. Enter your answer in simplified fraction form. ???
Still Need Help?
Join the QuestionCove community and study together with friends!
Sign Up
OpenStudy (anonymous):
Any clues?
OpenStudy (anonymous):
i guess my answer didn't help?
OpenStudy (anonymous):
no
OpenStudy (anonymous):
its 1/16
OpenStudy (anonymous):
right? 1/4 * 1/4= 1/16
Still Need Help?
Join the QuestionCove community and study together with friends!
Sign Up
OpenStudy (anonymous):
not the way i read the question no
OpenStudy (anonymous):
suppose there are four activities, lets call them A, B, C, D
what you wrote is the probability that both pick A
OpenStudy (anonymous):
how?
OpenStudy (anonymous):
is that clear?
OpenStudy (anonymous):
ooh so is it 1/16*1/4?
Still Need Help?
Join the QuestionCove community and study together with friends!
Sign Up
OpenStudy (anonymous):
or times 4 instead?
OpenStudy (anonymous):
no not divided by 4, times 4
OpenStudy (anonymous):
so the answer is 1/4?
OpenStudy (anonymous):
they both pick A or they both pick B or they both pick C or they both pick D
OpenStudy (anonymous):
yes, but there is an easier way to see this than that method
Still Need Help?
Join the QuestionCove community and study together with friends!
Sign Up
OpenStudy (anonymous):
how can I figure it out faster?
OpenStudy (anonymous):
imagine that the pick one after the other
the first person has to pick something
the second person has 4 choices, one of which is the same as the first person's pick
OpenStudy (anonymous):
so theres a 1/4 chance the second person will pick the same which is why it makes it 1/4?